### 29 Feb 2024 v3

# Dual-stage optimizer for systematic overestimation adjustment applied to multi-objective genetic algorithms for biomarker selection

Abstract
The challenge in biomarker discovery using machine learning from omics data lies in the abundance of molecular features but scarcity of samples. Most feature selection methods in machine learning require evaluating various sets of features (models) to determine the most effective combination. This process, typically conducted using a validation dataset, involves testing different feature sets to optimize the model’s performance. Evaluations have performance estimation error and when the selection involves many models the best ones are almost certainly overestimated. Biomarker identification with feature selection methods can be addressed as a multi-objective problem with trade-offs between predictive ability and parsimony in the number of features. Genetic algorithms are a popular tool for multi-objective optimization but they evolve numerous solutions thus are prone to overestimation. Methods have been proposed to reduce the overestimation after a model has already been selected in single-objective problems, but no algorithm existed capable of reducing the overestimation during the optimization, improving model selection, or applied in the more general multi-objective domain. We propose DOSA-MO, a novel multi-objective optimization wrapper algorithm that learns how the original estimation, its variance, and the feature set size of the solutions predict the overestimation. DOSA-MO adjusts the expectation of the performance during the optimization, improving the composition of the solution set. We verify that DOSA-MO improves the performance of a state-of-the-art genetic algorithm on left-out or external sample sets, when predicting cancer subtypes and/or patient overall survival, using three transcriptomics datasets for kidney and breast cancer.
Data and source code availability. The original gene expression data used in this study is from public repositories. The preprocessed data, source code, and detailed numerical results are available in a public server (github.com/UEFBiomedicalInformaticsLab/BIODAI/tree/main/DOSA_MO).
1 Introduction
Molecular biomarker discovery with machine learning (ML) is usually limited by data that includes many features but few samples [1]. This renders trained models prone to overfitting and the evaluation prone to estimation error. Hyperparameter tuning, including feature selection, is often a crucial aspect of the optimization process where the model’s performance is assessed using a training, validation, and test paradigm. This essential aspect involves choosing the most effective subset of features, essentially optimizing the ML’s model. Selecting the best model from many often leads to overestimation, given by a significant gap between validation set performance and actual, real-world performance, a phenomenon known as the “winner’s curse”. The models that fit the noise present in the validation set are advantaged, a phenomena sometimes referred as overfitting on the validation set [2]. This is also exacerbated by data scarcity. Seeking higher accuracies by expanding hyperparameter configurations can enhance model performance on validation sets. However, this often results in heightened overestimation, leading to reduced or potentially negative impact on test set performance.
In biomarker discovery, the focus is often on optimizing the accuracy of machine learning models using the selected molecular features, while also minimizing the number of features to ensure clinical feasibility and resource efficiency. Characterising all the best compromises (or trade-offs) between predictive value and feature set size is a multi-objective (MO) optimization problem [3, 4], and it can be solved by means of MO feature selection (MOFS) techniques [5, 6, 7, 8]. These techniques aim to identify not just a single best solution, as in single-objective (SO) problems, but rather a Pareto front of solutions. This front is the set of optimal solutions that illustrate the trade-offs between different objectives. However, all candidate solutions (or biomarker models selected by the employed MOFS technique) are evaluated on the validation set, which can result in the overestimation of the performance of the selected models.
K-fold cross-validation (CV) stands as the predominant methodology for machine learning assessment, with its advantages and limitations extensively explored, particularly in SO scenarios [9, 10, 11]. K-fold CV returns a model trained on all the available samples and an estimation of its performance computed by averaging $k$ CV results. The obtained model performance is usually underestimated when only one hyperparameter configuration is used. However, it tends to be overestimated when multiple configurations are evaluated, and the model (or feature set) with the best performance is returned [2]. This situation becomes particularly pronounced in MO approaches, where identifying multiple favorable trade-offs often leads to an increased number of models to be evaluated. A practical, illustrative example is presented in Cattelani et al. Fig. 1, where it is evaluated the performance of various gene expression-based molecular feature sets for classifying breast cancer subtypes [7]. In this MO setting, balanced accuracy and feature set size serve as evaluation metrics. The expected balanced accuracy tends to increase with the addition of more features. Similarly, overestimation – quantified as the difference between training and test set performance (the gap between “inner cv” and “test” results in the figures) – also grows as the estimation improves. These correlations may suggest that characteristics of the solutions might be predictive of the degree of overestimation.
Although various methods have been developed to enhance performance estimation in model selection using k-fold CV, their design and implementation have been limited to SO problems. Tsamardinos et al. [12] compared double CV, the Tibshirani and Tibshirani method [13], and nested CV in their ability to improve the estimation of the fitness for SO problems. These algorithms modify fitness estimations but do not change the model selection process; the chosen model remains the same as it would be using simple hyperparameter optimization with k-fold CV for model evaluation. Automated ML (AutoML) tools offer an approach designed to explore various model and hyperparameter combinations. These tools aim to identify and deliver the most effective model along with an assessment of its performance. In Tsamardinos et al. six AutoML tools were compared [14]. Of these, only one had a predictive performance estimation strategy that could adjusts for multiple model validations (limitedly to SO problems and not affecting model selection), while most of the tools have the necessity to withold a test set for an unbiased estimation of the performance of the winning model, thus loosing samples from the final model training.
No significant efforts have been directed toward improving performance estimation in MO problems, of which SO problems are a special case and which are more relevant and applicable to biomarker discovery in high-dimensional omics data. Moreover, while there are approaches for SO problems to enhance the performance estimation of a chosen model, these models are still selected based on unadjusted estimations. Consequently, this does not lead to any improvement in the actual model selection process. On the other hand, strategies that involve subtracting a constant value from model evaluation metrics, used for ranking multiple solutions, fall also short. This approach alone is insufficient to significantly impact the ranking order and thus remains ineffective for enhancing the model selection process. However, if the adjustment takes into account some characteristic of the models, such as the variance in performance during the validation phase or the feature set size, it might change their order. It may be argued that such an adjustment could improve the model selection if the model ranking would use the adjusted estimations. In summary, to the best of our knowledge, no previous work experimented the effectiveness of methods for mitigating the overestimation in MO problems using ML algorithms. Additionally, no previous work applied the adjustment to the performance estimation during the optimization of the solution set, thus affecting the selection.
Figure 1: Sequence of operations for external validation of DOSA-MO. A MO problem is defined with multiple objectives, e.g. cancer subtype classification, prediction of survival, and parsimony in the feature set size. A dataset (e.g. TCGA breast omics data) is fed to DOSA-MO for its optimization process in 3 steps. In step 1 it performs a k-fold CV with a wrapped MO optimizer (e.g. the NSGA3-CHS GA) and collects the solutions from all the folds. From each solution and objective a sample is constructed. It has the fitness expected by the wrapped MO, its SD, and the feature set size as independent variables, the overestimation (expected fitness minus fitness assessed on left-out set) as dependent variable, and the partial derivative on the HV with respect to this fitness measurement as sample weight. In step 2 these samples are used to train regression models for overestimation, and new adjusted objective functions are created. In step 3 a wrapped MO optimizer is run with the adjusted objective functions, impacting how the models are selected. The solution set is the output of DOSA-MO (e.g. a set of biomarkers). It might be beneficial to use a faster wrapped optimizer in step 1 than in step 3 since step 1 uses k-fold CV. An external dataset (e.g. SCAN-B) is used to externally validate the solution set.
Here we present the DOSA-MO (Dual-stage Optimizer for Systematic overestimation Adjustment in Multi-Objective problems), a cutting-edge algorithm aimed at predicting and adjusting for overestimation in MO problems. Initially, a wrapped MO optimizer is integrated with ML algorithms, to create a series of preliminary solutions (feature sets in our case study). These initial solutions are then used to train regression models that are specifically designed to predict performance overestimation. This prediction is based on characteristics of the solutions, such as the variance in evaluation metrics (for instance, the variance of balanced accuracy in internal validations) or the size of the feature set itself. As wrapped MO optimizer for our case study we used a modification of the Non-dominated Sorting Genetic Algorithm III (NSGA3) [15]: NSGA3 with Clone-Handling method and Specialized mutation operator (NSGA3-CHS, see Section 2.3). Subsequently, the algorithm executes a wrapped MO optimizer once more, but with a key distinction: it now utilizes the regression models to deliver adjusted fitness evaluations, enhancing the model selection process. The final resulting models are then trained on all the available samples, thus no samples are lost in order to compute their expected performance on new data. In our benchmarking study, which concentrates on selecting gene expression-based feature sets for cancer subtype classification and patient survival prediction, we have empirically demonstrated that DOSA-MO effectively mitigates overestimation and improves model selection. It consistently delivers improved performance estimations in biomarker discovery across diverse population-based cohort datasets. Additionally, two novel measures are introduced in our study: MO Performance Error (MOPE) and Pareto Delta ($P_{\Delta}$). To the best of our knowledge, these are the first to be designed to evaluate the discrepancy between the performance expected by the algorithm and the actual performance observed on new samples in MO problems. These assessments consider the entirety of the solution set determined by the MO optimizer, providing a comprehensive view of the algorithm’s effectiveness. The flow of a DOSA-MO use case is depicted in Fig. 1.
Disambiguations for the most technical terms can be found in Supplementary Section 3.
2 Methods
2.1 DOSA-MO: algorithm description
The DOSA-MO algorithm wraps a MO optimizer, serving two purposes: improve the estimate of the solutions’ performance, and increase the overall performance of the solution set (set of biomarker models in our case study). In our case study, these optimizers are represented by genetic algorithms (GAs) [16] specifically designed for MO problems and paired with supervised ML algorithms, such as the Gaussian Naïve Bayes (NB) classifier to distinguish cancer subtypes or the Cox Proportional-Hazards Model (Cox) for survival analysis. In more details, DOSA-MO consists of three steps.
1. Generating solution sets for overestimation prediction. It collects solutions to be used as training samples to learn how to adjust the objective functions that are used to evaluate solutions, such as the classification accuracy. This consists of running MO optimizers in a k-fold CV loop. For each fold, a solution set is produced using only training data, and its performance is measured on the left-out samples.
2. Training of regression models for overestimation. For each objective DOSA-MO trains a regression model on the samples collected during step 1. Each sample contains as independent variables three meta-features of the solutions that are potentially predictive of the overestimation. They are the original fitness (i.e. the fitness used by the optimizer to choose the best solutions, measured using only training data, applying inner k-fold CV in our case study), the standard deviation (SD) of that fitness (we measure it using bootstrap), and the number of features included in the solution (number of genes forming the biomarker in our case study). The dependent variable is the overestimation: the difference between the original fitness and the fitness computed on new data through CV. Solutions cannot be considered as equally important. We might expect solutions that are in crowded areas of the non-dominated front to be selected less often by a decision maker. Consequently, each sample is weighted according to the partial derivative of the HyperVolume (HV, constructed using the original fitnesses) [17] with respect to the considered solution and objective. The weights for each fold and objective are scaled to sum to 1. We minimize the absolute error, when allowed by the specific regression model, since the impact of an error on the HV is approximately linear for small errors. 3. Generating the solution set using adjusted performance. A second MO optimizer is deployed to generate a solution set (each solution refers to a feature set in our case study), with objectives that are adjusted by previously trained regression models for overestimation. Each original objective function is replaced by a pipeline that initially computes the function’s result, its SD, and the feature count of the solution. This data is then feed to the corresponding adjuster regression model, which predicts the overestimation. The final adjusted performance is calculated by deducting the predicted overestimation from the original fitness value. The solution set generated by the MO optimizer during this final run represents the output of the whole DOSA-MO.
Depending on the MO optimizers used, step 1 might be computationally expensive. A strategy to limit its cost when using GAs is described in Supplementary Section 1.3.
Step 2 includes computing the SD of the original fitness. Our method for doing that combining bootstrap with k-fold CV is described in Supplementary Section 1.4.
2.2 Pseudocode formulation of the DOSA-MO algorithm
In order to wrap any MO optimizers, the DOSA-MO must be polymorphic, so we define an abstract class MultiObjectiveOptimizer representing a generic MO optimizer (Pseudocode 1). It has a single method optimize that works with provided objectives and training data. The DOSA-MO is a MultiObjectiveOptimizer itself (Pseudocode 2).
class MultiObjectiveOptimizer
   method optimize(objectives, trainingData)
Pseudocode 1 MultiObjectiveOptimizer abstract class definition.
class DosaMO
      inherits MultiObjectiveOptimizer
   method new(
         tuningOptimizer, adjusterLearner,
         mainOptimizer):
      self.tuningOptimizer = tuningOptimizer
      self.adjusterLearner = adjusterLearner
      self.mainOptimizer = mainOptimizer
Ψ
   method optimize(objectives, trainingData):
      foldsData =
            createFolds(trainingData)
      foldHofs = [
            self.tuningOptimizer.optimize(
                  objectives, f.train)
            for f in foldsData]
      for i in 1::objectives.size:
         obj = objectives[i]
         weights =  [
            assignWeights(
            h.fitnessHyperboxes(), i)
            for h in foldHofs]
         adjuster = trainAdjuster(
               self.adjusterLearner, obj,
               foldHofs, foldsData, weights)
         adjustedObjectives[i] =
               createAdjustedObjective(
                     obj, adjuster)
      return self.mainOptimizer.optimize(
            adjustedObjectives, trainingData)
Pseudocode 2 DosaMO class definition.
The method new is just a simple constructor that saves the polymorphic parts of the algorithm: the tuningOptimizer, a MO optimizer used to create the samples for the adjusting regression, the actual regression model (adjusterLearner), and the mainOptimizer that is the MO optimizer that uses the adjusted objectives to produce the results for the user.
The DosaMO implementation of optimize first organizes the data into folds. The resulting object, foldsData, contains the data itself and also the description of how it is partitioned into folds. DosaMO then executes the tuningOptimizer on each fold and collects the results: a set of solutions for each fold. The set of solutions returned is optimizer-dependent in general, but in our experiments we used the non-dominated front of all the solutions that were explored. For each objective obj the algorithm assigns weights to the solutions: for each tuningOptimizer result set h, the solutions receive a weight that is proportional to the partial derivative of the HV of the belonging result set with respect to the solution and the current objective obj. An adjuster regression model is trained for the current objective obj using the function trainAdjuster that receives in input a regression model adjusterLearner, the current objective obj, the tuningOptimizer solution set for each fold, the data including folds information (foldsData), and the weigths of the samples (weights). The returned adjuster predicts how much the fitness of a solution changes between the training sample set and an unseen testing sample set. The function trainAdjuster has its own description below. Using the previous objective obj and the adjuster, a new adjusted objective is created that when evaluating a solution first uses the old fitness function to compute a temporary fitness and its SD, then adjustes this fitness by subtracting the prediction of the adjuster. Finally, DosaMO runs the mainOptimizer on the whole trainingData, using the adjusted objectives instead of the original objectives.
The trainAdjuster function trains a fitness adjuster regression model for one of the objectives using as samples the solutions resulting from running the tuningOptimizer on all the folds defined by foldsData. Each solution is assigned a weight proportional to the HV partial derivative with respect to the considered objective, with the weights for each fold that sum to 1.
trainAdjuster(
      adjusterLearner, obj, foldHofs,
      foldsData, weights):
   for i in 1::foldsData.size:
      allFitnesses =
            evaluateWithCV(
                   foldHofs[i], obj, foldsData[i])
      originalFitnesses[i] = allFitnesses.innerCV())
      stdDevs[i] = allFitnesses.innerCV_sd()
      testFitnesses[i] = allFitnesses.test()
      nFeatures[i] =
            [h.num_features() for h in foldHofs[i]]
   return adjusterLearner.fit(
         originalFitnesses, stdDevs, nFeatures,
         testFitnesses, weights)
Pseudocode 3 trainAdjuster function definition.
For each fold i as defined by foldsData, trainAdjuster prepares the samples for training the adjuster regressor (Pseudocode 3). The samples are prepared separately for each fold, then used together in training. The function evaluateWithCV assigns two fitnesses to each solution: one previously computed on the train data of the current fold i (using a nested k-fold CV in our case study), and another computed on the test data. It also computes the SD of the train fitness with the bootstrap method (Supplementary Section 1.4). The differences between train and test performance are the values that the regression will learn to predict. The regression has 3 input meta-features: the original fitness, i.e. the performance on the train data, the SD of the original fitness, and the number of the features that are included in the solution (number of genes in our case study).
2.3 Case study overview
We have benchmarked DOSA-MO in the context of MOFS for cancer biomarker discovery. Our goal was to identify biomarkers for classifying cancer subtypes and survival prediction in kidney and breast cancer patients. This was done using gene-based transcriptomic datasets from The Cancer Genome Atlas (TCGA) project [18]. For breast cancer, an additional cohort from The Sweden Cancerome Analysis Network - Breast (SCAN-B) [19] serves as an external validation set. A description of the datasets and their preprocessing is in Supplementary Section 1.5. As wrapped MO optimizer, we used a novel modification of NSGA3 [15]: NSGA3-CHS, with NB or Support Vector Machine (SVM) as inner classifiers, and Cox as inner survival model. We used both internal k-fold CV and external validation to compare the unadjusted optimizer (abbreviated as “zero”) with adjustments by 5 different regression models: weighted median (dummy), pruned decision tree (ptree), random forest (RFReg), support vector regression (SVR), and SVR with optimized regularization parameters (rSVR). The regression models are described in detail in Supplementary Section 1.2.
Datasets Objectives Inner models Validation
Kidney cancer Subtype classification (balanced accuracy), set size (root-leanness) NB CV
Kidney cancer Overall survival (c-index), set size (root-leanness) Cox CV
Kidney cancer Subtype classification (balanced accuracy), overall survival (c-index), set size (root-leanness) NB/Cox CV
Breast cancer Subtype classification (balanced accuracy), set size (root-leanness) NB, SVM CV
Breast cancer Subtype classification (balanced accuracy), set size (root-leanness) NB, SVM External (SCAN-B)
Breast cancer Overall survival (c-index), set size (root-leanness) Cox CV
Table 1: Overview of datasets, objectives, inner models, and validation methods used in the analysis of kidney and breast cancer.
NSGA3-CHS is an instantiation of the more general algorithm NSGA*. They are both defined in Supplementary Section 1.1. The considered experimental setups are listed in Table 1 and described in detail in Supplementary Section 1.6, where it is also defined the root-leanness, used as fitness function for the parsimony of the feature sets. Classification and survival prediction performance are measured respectively with balanced accuracy and concordance index (c-index).
2.4 Measuring overestimation in multi-objective problems
We propose two new methods to measure the error of the performance estimates for the solutions to MO problems: MOPE and $P_{\Delta}$. The data-driven approach complicates the measurements, as CV yields varied performance metrics. These include the performance anticipated by the optimizer, based on training data, and the performance measured on left-out/new samples post-optimization. To our knowledge, there is no established metric for the error in evaluating the predictive performance of solutions in MO problems within a CV framework. In SO scenarios, one might simply assess the absolute difference between the fitness expected from training data and the fitness observed on new data. However, in MO scenarios, it is crucial to consider each solution’s contribution to the Pareto front. We introduce two novel metrics to measure this estimation error in MO CV setups.
2.4.1 Multi-objective performance error (MOPE)
We define the MOPE $E_{\upsilon}$ starting from the HV [17] computed on the train performance ($H_{\iota}$) and the Cross HyperVolume (CHV) ($H_{\upsilon}$). $H_{\iota}$ and $H_{\upsilon}$ have been formally defined by Cattelani et al. [8]. Since $H_{\upsilon}$ is a family of functions, with the specific instantiation that depends on the user provided function $\upsilon$, $E_{\upsilon}$ is a family of functions too.
$E_{\upsilon}(X,X^{\prime})=\left\lvert H_{\iota}(X,X^{\prime})-H_{\upsilon}(X,X^{\prime})\right\rvert$ (1)
Where $X$ encodes the performance of the solutions evaluated on the train data, and $X^{\prime}$ the performance evaluated on the test data. In our case study we use the same instantiation of $\upsilon$ as in Cattelani et al.: the function $\lambda$ [8].
$E_{\upsilon}$ has a simple definition and can be computed very efficiently if the experimental setup already includes the measuring of $H_{\iota}$ and $H_{\upsilon}$. It can be seen intuitively as the difference between the aggregated performance of all the solution set as expected by the optimizer taking into account only train data, and the aggregated performance of the same solutions when applied on never before seen test data by decision makers that must choose their preferred solution informed by train performance only.
2.4.2 Pareto delta ($P_{\Delta}$)
The MOPE, derived directly from the training HV and CHV, does not specify the exact sources of discrepancies between these measures. Although CHV accounts for differences between training and testing performance, variations in solution performance of opposite sign can offset each other, resulting in a lower MOPE despite significant train-test discrepancies. To tackle this issue, we introduce a supplementary metric, the $P_{\Delta}$. It has the property of being equal to 00 only when there is no difference in performance between train and test data for all elements of the solution set. To calculate the $P_{\Delta}$, we sum up the absolute error in fitness estimation for each solution, multiplying it by the derivative of the HV concerning that specific objective and solution. Then, we compute the average across all objectives.
Let $n$ be the number of solutions and $m$ the number of objectives. Let $X$ be an $n\times m$ matrix where $x_{i,j}$ is the train performance for the $j^{th}$ objective of the $i^{th}$ solution, with $1\leq i\leq n$ and $1\leq j\leq m$. Similarly, let $X^{\prime}$ be an $n\times m$ matrix where $x^{\prime}_{i,j}$ is the test performance for the $j^{th}$ objective of the $i^{th}$ solution. Let $H_{\iota}$ be the HV computed on the train performance $X$. $\partial H_{\iota}/{\partial x_{i,j}}$ is the partial derivative of the HV with respect to $x_{i,j}$. We define the $P_{\Delta}$ function as
$P_{\Delta}(X,X^{\prime})=\frac{1}{m}\sum_{j=1}^{m}\sum_{i=1}^{n}\left\lvert x_{i,j}-x^{\prime}_{i,j}\right\rvert\frac{\partial H_{\iota}}{\partial x_{i,j}}$ (2)
In the special cases with 00 dimensions or 00 solutions we define $P_{\Delta}$ as 00. As long as the differences $\left\lvert x_{i,j}-x^{\prime}_{i,j}\right\rvert$ are small, $P_{\Delta}$ is proportional to the difference between the volume of the geometrical union of the space under the train and test fronts and the volume of the intersection of the same two spaces.
3 Results
(a) (b) (c)
Figure 2: Scatter plots depicting solutions from external validation on breast cancer transcriptomics data using SVM as inner model. MO optimization of balanced accuracy for subtypes prediction and root-leanness. Horizontally, the number of features is depicted for simplicity. For each solution it is shown the performance measured in the inner CV, i.e. the performance expected by the optimizer, the performance of the model trained on the TCGA breast set and tested on the same set, and the performance of the same model on the external SCAN-B set. The lines are interpolating splines. (a) Using the unadjusted optimizer. (b) Using SVR as regression model for fitness adjustment. (c) Using RFReg as regression model for fitness adjustment.
We compare 6 different instantiations of the DOSA-MO by varying the regression model used for the adjustment, including the zero regression model, equivalent to not applying any adjustment. The regression models receive as input 3 meta-features of the solutions correlated with the overestimation: the original fitness, its SD, and the feature set size (number of genes). For an example of correlations between meta-features and overestimation see Supplementary Section 2.1. We use NSGA3-CHS as wrapped model. The tests are repeated with 8 different combinations of validation type, datasets, objectives, and classification inner model. For each combination we report two measures for the accuracy of the fitness estimation: the MOPE (Section 3.1.1) and the $P_{\Delta}$ (Section 3.1.2). Additionally, we report a measure of overall performance of the optimizers: the CHV (Section 3.2). A focus on the best solution sets from the external validation is shown in Fig. 2 and discussed in Supplementary Section 2.2.
3.1 Overestimation in feature selection for biomarker discovery
We measured the error of the performance estimates with two methods: the MOPE and the $P_{\Delta}$. The MOPE quantifies the discrepancy between the HV computed on the train performance and the CHV. In contrast, the $P_{\Delta}$ captures differences between expected performance and actual performance on new samples, focusing on the more granular level of individual solutions.
3.1.1 Comparative analysis using multi-objective performance error
(a) (b) (c)
(d) (e) (f)
Figure 3: MOPE results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification, and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The MOPE evaluates the DOSA-MO algorithm’s capability in reducing the performance estimation error (Section 2.4.1). The MOPE outcomes from various regression-based adjusters are reported in Fig. 3.
Significantly, ptree and RFReg often outperform other overestimation predictors. Therefore, for users prioritizing the reduction of MOPE, ptree and RFReg regressors emerge as particularly effective choices. SVR-based estimators of overfitting appear to be the least effective. ptree and RFReg regressors perform particularly well in external validation scenarios, where an external dataset (SCAN-B) is used to evaluate the models (biomarker sets), that were selected according to their adjusted performance learned from the initial TCGA breast cohort.
3.1.2 Comparative analysis using the Pareto delta
(a) (b) (c)
(d) (e) (f)
Figure 4: $P_{\Delta}$ results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The $P_{\Delta}$ measure (Section 2.4.2) considers the difference between a model’s predicted and actual test performance for each individual solution. This makes it particularly relevant and more informative than the MOPE in the common situation where users are presented with multiple solutions but will ultimately select only one. The $P_{\Delta}$ across the experimental setups is shown in Fig. 4.
ptree and RFReg regressors consistently outshine others when reducing the $P_{\Delta}$. Interestingly, while no significant differences are noted in MOPE between the dummy model and the unadjusted optimizer, all kidney cancer setups with proper regressors have a lower $P_{\Delta}$ compared to the unadjusted optimizer. rSVR performs better than the zero regressor in all scenarios except one. RFReg and ptree always yield lower $P_{\Delta}$s than the zero regressor, including in breast cancer setups, highlighting their promise in both overestimation measures.
3.2 Evaluating DOSA-MO’s effectiveness in MO feature selection
(a) (b) (c)
(d) (e) (f)
Figure 5: CHV results for internal k-fold CV (a-e) and external validation (f). (a) Kidney cancer, subtype classification and root-leanness. (b) Kidney cancer, overall survival prediction and root-leanness. (c) Kidney cancer, overall survival prediction, subtype classification and root-leanness. (d) Breast cancer, subtype classification and root-leanness. (e) Breast cancer, overall survival prediction and root-leanness. (f) External validation for breast cancer, subtype classification and root-leanness. Error bars represent SD between folds.
The previous results highlight DOSA-MO’s success in reducing performance estimation errors in ML-driven feature selection for biomarker discovery. It is also important to evaluate if DOSA-MO enhances the quality of the produced feature sets: the final output for biomarker selection models considered for clinical validation. To assess the impact on the overall feature selection process, we calculated the CHV for each experimental setup (see Fig. 5).
Tree-based regressors reduce estimation errors but do not significantly improve the quality of solution sets generated by DOSA-MO. This indicates that using these regressors for overestimation prediction may not always enhance solution sets, despite reducing quality estimation errors. Interestingly, dummy and SVR-based MO optimizers consistently outperform the CHV of unadjusted models, highlighting the effectiveness of overestimation prediction for superior feature selection outcomes. Specifically, SVR yelds the best CHV in all setups except in scenarios involving the overall survival objective, while the dummy model excels in setups that include survival prediction.
4 Discussion
We introduced DOSA-MO, a novel wrapper MO optimizer designed to adjust performance measures for overestimation in ML-driven MO problems, improving the solution sets that are produced. This approach has been successfully validated across multiple transcriptomic-based datasets commonly used for biomarker discovery in cancer subtype classification and survival prediction. Our benchmarking extends to a scenario where biomarker models, trained in one population-based cohort (such as TCGA), are applied to classify cancer subtype in a second cohort (such as SCAN-B). This method effectively enables biomarker discovery and validation across multiple cohorts. Additionally, we have developed two innovative measures for evaluating the performance of MO algorithms: the MOPE and the $P_{\Delta}$. According to both metrics, the DOSA-MO algorithm demonstrated improved performance estimates in all tested cases, particularly when using decision tree-based regression models for predicting overestimation.
Our study found that even a basic regression model, which learns solely the weighted average of overestimation and is used for fitness adjustment, resulted in improved overall performance compared to the unadjusted optimizer in 7 out of 8 experimental setups, as indicated by the CHV metric. Likewise, the MO optimizer directed by the SVR model, which predicts overestimations, surpassed the unadjusted optimizer in 7 out of 8 setups, affirming its effectiveness. Notably, the MO optimizer guided by the simpler dummy regression model excelled in the three setups involving survival prediction, while the SVR model proved superior in the five setups focused on subtype classification and feature set size. The rSVR regression model, that selects the hyperparameters for the regularization strength of the SVR with random search, utilized for predicting overestimation, underperformed with respect to dummy and SVR regression models. A regularization that works well while cross-validating on solutions obtained by an unadjusted optimizer, is not as effective when the GA runs with a bigger population, more generations, and the adjustment is applied to the evaluation of the individuals.
Predicting the overestimation is more complex when the adjustment to the fitness is applied during the optimization as in our approach. The fitness adjustments alters the optimizer’s exploration path, leading to a divergence in the solution distribution from the one used to train the regressors for overestimations, even within their applicability domain. This creates a “moving target” problem, commonly addressed in other contexts, like artificial neural networks, through incremental optimization across multiple epochs. Although executing more computationally intensive MO GAs is possible, further research is essential to enhance our understanding and address the challenges of the applicability domain and the moving target problem in these contexts.
DOSA-MO is versatile and can be applied to any ML problem, including the ones requiring a MO evaluation framework, making it particularly well-suited for biomarker discovery in high dimensional molecular datasets. It has been experimentally validated with cancer transcriptomics data, demonstrating its applicability and effectiveness in this domain.
Funding
This work was supported by the Academy of Finland [grant agreements 336275, 332510, and 358037], the Jane and Aatos Erkko Foundation [210026], and the Sigrid Jusélius Foundation.
Acknowledgments
The computational analyses were performed on servers provided by the UEF Bioinformatics Center, University of Eastern Finland, Finland.